Determine the effective capacitance of the following combination of capacitors given that the capacitance of each capacitor is 3 μF.
1
Expert's answer
2019-03-28T06:04:55-0400


And so, the first step

Сeq1 = C1 + C2

Сeq1 = 3 μF + 3 μF = 6 μF

The second step

1/Ceq = 1/Ceq1 + 1/C3

Ceq = (Ceq1 x C3) / (Ceq1 + C3)

Ceq = (6 μF x 3 μF) / (6 μF + 3 μF) = 2 μF

Answer: 2 μF